# A LoD of Gaussians: Unified Training and Rendering for Ultra-Large Scale Reconstruction with External Memory

### v2

@@ Abstract.
Gaussian Splatting has emerged as a high-performance technique for novel view synthesis, enabling real-time rendering and high-quality reconstruction of small scenes. However, scaling to larger environments has so far relied on partitioning the scene into chunks—a strategy that introduces artifacts at chunk boundaries, complicates training across varying scales, and is poorly suited to unstructured scenarios such as city-scale flyovers combined with street-level views. Moreover, rendering remains fundamentally limited by GPU memory, as all visible chunks must reside in VRAM simultaneously. We introduce A LoD of Gaussians, a framework for training and rendering ultra-large-scale Gaussian scenes on a single consumer-grade GPU—without partitioning. Our method stores the full scene out-of-core (e.g., in CPU memory) and trains a Level-of-Detail (LoD) representation directly, dynamically streaming only the relevant Gaussians. A hybrid data structure combining Gaussian hierarchies with Sequential Point Trees enables efficient, view-dependent LoD selection, while a lightweight caching and view scheduling system exploits temporal coherence to support real-time streaming and rendering. Together, these innovations enable seamless multi-scale reconstruction and interactive visualization of complex scenes—from broad aerial views to fine-grained ground-level details.
Level of Detail, Gaussian Splatting, Large-Scale Reconstruction
isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Rasterizationccs: Computing methodologies Visibilityccs: Computing methodologies Machine learning
Figure 1. We introduce a fully hierarchical 3D Gaussian representation trained directly across unstructured, multi-scale image sets—including street-level and aerial views—without scene partitioning. Our method maintains a consistent global scene model, eliminating boundary artifacts typical of chunked approaches. A hybrid Level-of-Detail system combines Gaussian hierarchies with Sequential Point Trees, enabling dynamic, view-dependent streaming and LoD selection. The entire model resides in external memory, with only a small, adaptive subset loaded on demand, allowing real-time rendering and training of 60M+ Gaussians on consumer GPUs ($\leq 24GB$ VRAM). This is the first approach to achieve bounded-memory training and rendering of city-scale Gaussian splats at interactive rates.
[…]
Visual Abstract.
1. Introduction
@@
Given a set of posed images of a 3D scene, the task of novel view synthesis (NVS) is to generate plausible images of the scene from unseen viewpoints. Early approaches achieved this via image-based blending (Zhang et al., 2019), but the introduction of Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) marked a breakthrough, enabling high-quality results by optimizing an implicit volumetric scene representation through a multi-layer perceptron. More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) extended this paradigm to an explicit representation: a set of Gaussian primitives that are efficiently rasterized using splatting techniques (Zwicker et al., 2001), replacing costly ray marching and allowing real-time rendering with fast convergence.
@@
Despite these advances, both NeRF and 3DGS remain constrained by memory bottlenecks when applied to large-scale environments. Prior methods address this by dividing scenes into smaller chunks (Xu et al., 2023; Tancik et al., 2022; Li et al., 2024; Kerbl et al., 2024; Liu et al., 2024a; Chen et al., 2024b; Lin et al., 2024), training each independently before merging results. While chunking strategies mitigate memory usage during training, they introduces several key limitations:
  1. (1)
@@
    View-chunk misalignment: Camera views often span multiple chunks, especially in open environments or multi-scale datasets (e.g., combining aerial and street-level images). This makes chunk boundaries arbitrary and hard to define, complicating scene partitioning and training.
  2. (2)
@@
The simplest and most robust alternative to chunking is to avoid splitting altogether. With A LoD of Gaussians, we introduce a seamless pipeline that enables training and rendering of ultra-large-scale scenes directly—on a single consumer-grade GPU—without any form of scene partitioning. To handle scenes that exceed available VRAM, we store all Gaussian data in CPU RAM and dynamically stream only those visible from the current training view into GPU memory. However, a single far-away view could still require access to the full scene. To address this, we construct a hierarchical Level-of-Detail (LoD) model inspired by Kerbl et al. (2024), loading detail proportional to view distance. Maintaining this hierarchy during training is non-trivial, as Gaussian properties evolve dynamically. We propose a novel hierarchy densification strategy, adapted from the MCMC-style spawning (Kheradmand et al., 2024), to support stable, progressive refinement. Efficient view-dependent selection from the hierarchy is challenging for large models. Instead of full tree traversal, we adopt Sequential Point Trees (SPTs) (Dachsbacher et al., 2003), originally developed for point cloud LoD rendering. Our Hierarichal SPT version allows us to compute the correct LoD cut efficiently for rendering individual views and camera paths. Finally, to reduce CPU-GPU data transfer overhead, we introduce a lightweight caching system that tracks recently used Gaussians and reuses them across training iterations. In summary, we make the following contributions:
  1. (1)
    We propose a novel hierarchy densification strategy that enables dynamic expansion and restructuring of Gaussian representations during training.
  2. (2)
    We leverage the Sequential Point Tree (SPT) data structure and adapt it for large-scale Gaussian Splatting to perform fast, parallelizable LoD cuts for efficient training and real-time rendering.
  3. (3)
@@
2. Related Work
@@
Reconstructing large-scale scenes from images has long been a central challenge in visual computing. Traditional approaches relied on Structure-from-Motion pipelines, such as Building Rome in a Day (Agarwal et al., 2009) and COLMAP (Schönberger and Frahm, 2016), to recover geometry from unordered photo collections. Differentiable rendering techniques, notably NeRF (Mildenhall et al., 2020) and 3DGS (Kerbl et al., 2023), marked a paradigm shift by optimizing volumetric scene representations. Extensions of NeRF to large scenes typically employ scene partitioning (Tancik et al., 2022; Xu et al., 2023) or multi-GPU training strategies (Li et al., 2024). Similarly, most large-scale 3DGS pipelines adopt chunk-based training: Hierarchical 3DGS (Kerbl et al., 2024) trains chunks independently and then merges them into a global LoD hierarchy; CityGaussian (Liu et al., 2024a) combines chunked training with per-chunk LoD selection using LightGaussian (Fan et al., 2024); and VastGaussian (Lin et al., 2024) introduces decoupled appearance modeling and progressive partitioning. Horizon-GS (Jiang et al., 2024) integrates divide-and-conquer strategies with LoD mechanisms from (Ren et al., 2024), specifically targeting hybrid aerial/street-view datasets. Grendel (Zhao et al., 2024) avoids spatial chunking by distributing training images across GPUs, such that each device renders a disjoint screen region. Another research direction focuses on extracting geometric proxies from large-scale 3DGS scenes (Li et al., 2025; Liu et al., 2025; Chen et al., 2024b). These methods leverage TSDF fusion and geometric losses to generate multiple meshes, which are fused and rendered efficiently using traditional rasterization.
Compression
Lowering memory consumption is essential to scaling 3DGS to larger scenes. Compression techniques typically fall into three overlapping categories: compaction, which reduces the number of Gaussians; attribute compression, which reduces the storage per primitive; and structured representations, which organize data to improve compressibility. Pure compaction strategies include (Ren et al., 2024; Kim et al., 2024; Fang and Wang, 2024; Mallick et al., 2024; Cheng et al., 2024; Liu et al., 2024c); hybrid methods that also compress attributes include (Wang et al., 2024b; Papantonakis et al., 2024a; Fan et al., 2024; Girish et al., 2024); and pure attribute compression is addressed by (Niedermayr et al., 2024). Methods combining all three dimensions include (Chen et al., 2024a; Lee et al., 2024), while (Wang et al., 2024a; Ye et al., 2025; Navaneet et al., 2023; Liu et al., 2024b) emphasize attribute compression with structural priors. Finally, Lu et al. (2023) focus on hierarchical anchoring. These techniques are largely orthogonal to our work—many are applied post hoc or modify densification policies—and could be combined with our pipeline for further gains in memory and performance.
@@
Level-of-detail techniques reduce the complexity of distant scene content to accelerate rendering. In the context of 3DGS, LoD approaches have been explored for enabling efficient rendering on memory-constrained or mobile devices. Compression-based strategies include attribute quantization via codebooks, pruning of low-impact Gaussians, and adapting the degree of spherical harmonics per primitive (Papantonakis et al., 2024b; Niedermayr et al., 2024; Fan et al., 2024; Fang and Wang, 2024; Huang et al., 2025; Niemeyer et al., 2025; Seo et al., 2024). Scaffold-GS (Lu et al., 2023) introduced learned feature vectors anchored to reference Gaussians, with an MLP generating associated Gaussians at render time. This concept was extended to hierarchical LoD rendering in Octree-GS (Ren et al., 2024), enabling real-time control over detail levels through spatial subdivision.
3. Preliminaries
We briefly review 3D Gaussian Splatting and describe its memory usage characteristics. We then introduce the key data structures we used: Gaussian hierarchies and Sequential Point Trees.
3.1. 3D Gaussian Splatting
3D Gaussian Splatting (Kerbl et al., 2023) models a radiance field using a set of spatially distributed Gaussians, each with mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, RGB base colors $\mathbf{b}_{i}\in\mathbb{R}^{3}$ and covariance matrices $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$. The covariance is parameterized via a diagonal scaling matrix $\mathbf{S}_{i}=\text{diag}(s^{1}_{i},s^{2}_{i},s^{3}_{i})$ and an orthonormal rotation matrix $\mathbf{R}_{i}$:
$\boldsymbol{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}.$
Each Gaussian also stores an opacity $\sigma_{i}$ and view-dependent color, modeled using spherical harmonics (SH) coefficients $\mathbf{f}_{i}^{d}$. The SH degree $d$ controls expressiveness, with each Gaussian requiring $\sum_{j=1}^{d}3\cdot(2j+1)$ parameters. For rendering, all $N$ Gaussians are sorted by distance to the camera and a discrete approximation of the volume rendering equation is evaluated for every pixel $\mathbf{x}$ with corresponding view direction $\mathbf{v}$:
$\mathbf{C}(\mathbf{x})=\sum_{i=1}^{N}\mathbf{c}_{i}(\mathbf{v})\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$
@@ 3. Preliminaries
$\alpha_{j}(\mathbf{x})=\sigma_{j}e^{-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu^{\prime}}_{j})\boldsymbol{\Sigma^{\prime}}(\mathbf{x}-\boldsymbol{\mu^{\prime}}_{j})^{T}}.$
@@
Here $\boldsymbol{\mu^{\prime}}$ and $\boldsymbol{\Sigma^{\prime}}$ denote the projected 2D mean and covariance on the image plane, obtained by applying an affine approximation of the projective transform (Zwicker et al., 2001). The rendering pipeline is fully differentiable, enabling end-to-end optimization with ADAM using a combination of $\mathcal{L}_{1}$ and SSIM losses.
3.2. Memory in 3DGS
@@
Standard 3DGS pipelines store the full set of per-Gaussian attributes, training images, and optimizer state in GPU memory (VRAM). Figure 15 summarises the amount of memory that the various components of a single Gaussian require. Additional temporary allocations occur during forward and backward passes (e.g., for sorting and gradient accumulation). This overhead varies with the scale of the Gaussians and effectiveness of culling strategies, but can be roughly upper-bounded by around 800 bytes per Gaussian in practice. This limits training to roughly $500$K Gaussians per GB of VRAM, imposing strict constraints on the detail and extent of reconstructions.
3.3. Gaussian Hierarchies and Sequential Point Trees
@@
Gaussian hierarchies, introduced in (Kerbl et al., 2024), recursively merge nearby Gaussians into a tree, where each non-leaf node approximates its children, and leaves correspond to the original Gaussians. A cut is defined by a condition $c_{\text{hier}}(i,\text{cam})$ evaluated in a breadth-first search (BFS). If a node satisfies the cut condition, it is added to the cut set and its children are skipped; otherwise, the BFS continues. A proper cut set includes no parents or children of any included node and thus provides a view-adaptive LoD representation.
The cut condition used in (Kerbl et al., 2024) is:
$c_{\text{hier}}(i,\text{cam})=\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i),\quad m_{d}(i)=\frac{T}{\max_{j}s_{i}^{j}},$
@@
where $\mathbf{p}_{cam}$ is the camera position, $T$ is a global LoD threshold, and $m_{d}(i)$ is the minimum acceptable distance for viewing Gaussian $i$. This ensures that if $i$ is in the cut set, its parent failed the condition:
$m_{d}(\text{parent}(i))>\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
Sequential Point Trees (SPTs) (Dachsbacher et al., 2003), originally developed for point cloud LoD rendering, can be adapted for Gaussians. They enforce a more constrained cut condition:
$c_{\text{SPT}}(i,\text{cam})=m_{d}(\text{parent}(i))>\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
This condition is evaluated for all Gaussians in parallel, using the shared root-camera distance $\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$. It requires storing only sorted pairs $\big{(}m_{d}(i),m_{d}(\text{parent}(i))\big{)}$, significantly reducing memory compared to full Gaussian hierarchies. Note that cuts are guaranteed to be proper, with nodes where $m_{d}(i)<m_{d}(\text{parent}(i))$ never being selected for the cut. To optimize cuts, Gaussians are sorted by $m_{d}(\text{parent}(i))$ in descending order. A binary search determines the cutoff index $N$, above which Gaussians are too fine to be rendered. The downside of the approach is that all Gaussians in the same SPT share the same level of detail dictated by the distance from the camera to the root node. This can lead to Gaussians with $m_{d}(i)>\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}$ being rendered, even though they would be too coarse for the current view. To counteract this issue, we define:
$M_{d}(i)=m_{d}(i)+\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2},$
as a conservative minimum distance function. By the triangle inequality, selecting Gaussians satisfying $M_{d}(i)\leq\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$ guarantees $m_{d}(i)\leq\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}$. In turn, this means that Gaussians that are further away from the camera than the root node will be selected at a higher level of detail. SPTs are best suited for tightly grouped Gaussians observed from distances greater than their mutual spacing. Their compact memory footprint and parallel evaluation make them well-suited for large-scale scenes. Figure 2 visualizes both hierarchy types and their LoD cuts for a toy example.
Figure 2. A Sequential Point Tree and Gaussian hierarchy represent the same 5 Gaussians at varying levels of detail shown with four possible hierarchy cuts in red. Horizontal lines in the SPT show the binary search result and horizontal lines the distance cut.
[…]
[SPT vs Gaussian hierarchy]Sequential Point Tree and Gaussian hierarchy datastructure for a set of 9 Gaussians along with various levels of detail and their cuts
4. Method

### v3

@@ Abstract
Gaussian Splatting has emerged as a high-performance technique for novel view synthesis, enabling real-time rendering and high-quality reconstruction of small scenes. However, scaling to larger environments has so far relied on partitioning the scene into chunks—a strategy that introduces artifacts at chunk boundaries, complicates training across varying scales, and is poorly suited to unstructured scenarios such as city-scale flyovers combined with street-level views. Moreover, rendering remains fundamentally limited by GPU memory, as all visible chunks must reside in VRAM simultaneously. We introduce A LoD of Gaussians, a framework for training and rendering ultra-large-scale Gaussian scenes on a single consumer-grade GPU without partitioning. Our method stores the full scene out-of-core (e.g., in CPU memory) and trains a Level-of-Detail (LoD) representation directly, dynamically streaming only the relevant Gaussians. A hybrid data structure combining Gaussian hierarchies with Sequential Point Trees enables efficient, view-dependent LoD selection, while a lightweight caching and view scheduling system exploits temporal coherence to minimize the loading overhead. Together, these innovations enable seamless multi-scale reconstruction and interactive visualization of complex scenes—from broad aerial views to fine-grained ground-level details.
1 Introduction
@@
Given a set of posed images of a 3D scene, the task of novel view synthesis (NVS) is to generate plausible images of the scene from unseen viewpoints. Early approaches achieved this via image-based blending (DeepBlending), but the introduction of Neural Radiance Fields (NeRF) (NeRF) marked a breakthrough, enabling high-quality results by optimizing an implicit volumetric scene representation through a multi-layer perceptron. More recently, 3D Gaussian Splatting (3DGS) (3DGS) extended this paradigm to an explicit representation: a set of Gaussian primitives that are efficiently rasterized using splatting techniques (zwicker2001EWA), replacing costly ray marching and allowing real-time rendering with fast convergence.
@@
Despite these advances, both NeRF and 3DGS remain constrained by memory bottlenecks when applied to large-scale environments. Prior methods address this by dividing scenes into smaller chunks (GridNeRF; BlockNeRF; Hierarchical; CityGaussian; GigaGS; VastGaussian), training each independently before merging results. While chunking strategies mitigate memory usage during training, they introduce several key limitations:
  1. 1.
@@
    View-chunk misalignment: Camera views often span multiple chunks, especially in open environments or multi-scale datasets (e.g., combining aerial and street-level images). This makes chunk boundaries arbitrary and hard to define, complicating scene partitioning and training. We identify two artifacts caused by view-chunk misalignment: Chunk bleeding occurs when Gaussians extend out of their current chunk and obscure neighbouring chunks after merging. When an occluder on a training image is not part of the current chunk, its ”ghostly” outline will be trained into the wrong chunk, leading to Chunk ghosting. Examples of these artifacts can be seen in Figures6 and 11.
  2. 2.
@@
The simplest and most robust alternative to chunking is to avoid splitting altogether. With A LoD of Gaussians, we introduce a seamless pipeline that enables training and rendering of ultra-large-scale scenes directly on a single consumer-grade GPU, without any form of scene partitioning (see Figure 7). To handle scenes that exceed available VRAM, we store all Gaussian data in CPU RAM and dynamically stream only those visible from the current training view into GPU memory. Still, a single distant view could require access to the full scene. To address this, we construct a hierarchical Level-of-Detail (LoD) model inspired by Hierarchical, loading detail proportional to view distance. Maintaining this hierarchy during training is non-trivial, as Gaussian properties evolve dynamically. We propose a novel hierarchy densification strategy, adapted from the MCMC-style spawning (3DGSasMCMC), to support stable, progressive refinement. Efficient view-dependent selection from the hierarchy is challenging for large models. Instead of full tree traversal, we adopt Sequential Point Trees (SPTs) (SPT), originally developed for point cloud LoD rendering. Our Hierarchical SPT version allows us to compute the correct LoD cut efficiently for rendering individual views and camera paths. Finally, to reduce CPU-GPU data transfer overhead, we introduce a lightweight caching system that tracks recently used Gaussians and reuses them across training iterations. In summary, we make the following contributions:
  1. 1.
    We propose a novel hierarchy densification strategy that enables dynamic expansion and restructuring of Gaussian LoD representations during training.
  2. 2.
    We leverage the Sequential Point Tree (SPT) data structure and adapt it for large-scale Gaussian Splatting to perform fast, parallelizable LoD cuts for efficient training and rendering.
  3. 3.
@@
  5. 5.
    We verify trough extensive evaluations that the pipeline succeeds in reconstructing huge datasets with complex camera distribution, where SOTA divide-and-conquer methods fail even with larger VRAM budgets.
2 Related Work
@@
Reconstructing large-scale scenes from images has long been a central challenge in visual computing. Traditional approaches relied on Structure-from-Motion (SfM) pipelines to recover geometry from unordered photo collections (RomeInADay; COLMAP). Differentiable rendering techniques, notably NeRF (NeRF) and 3DGS (3DGS), marked a paradigm shift by optimizing volumetric scene representations. Extensions of NeRF to large scenes typically employ scene partitioning (BlockNeRF; GridNeRF) or multi-GPU training strategies (NerfXL). Similarly, most large-scale 3DGS pipelines adopt chunk-based training: Hierarchical-3DGS (H-3DGS) (Hierarchical) trains chunks independently and then merges them into a global LoD hierarchy; CityGaussian (CityGaussian) combines chunked training with per-chunk LoD selection using LightGaussian (LightGaussian); and VastGaussian (VastGaussian) introduces decoupled appearance modeling and progressive partitioning. Horizon-GS (HorizonGS) integrates divide-and-conquer strategies with LoD mechanisms from OctreeGS, specifically targeting hybrid aerial/street-view datasets. GrendelGS (Grendel) avoids spatial chunking by distributing training images across GPUs, such that each device renders a disjoint screen region. Another research direction focuses on extracting geometric proxies from large-scale 3DGS scenes (ULSRGS; CityGaussianV2; GigaGS). These methods leverage TSDF fusion and geometric losses to generate multiple meshes, which are fused and rendered efficiently using traditional rasterization.
@@
Level-of-detail techniques reduce the geometric complexity of distant scene content to accelerate rendering. In the context of 3DGS, LoD approaches have been explored for enabling efficient rendering on memory-constrained or mobile devices. Compression-based strategies include attribute quantization via codebooks, pruning of low-impact Gaussians, and adapting the degree of spherical harmonics per primitive (ReducingTheMemoryFootprint; Compressed3DGaussianSplatting; LightGaussian; MiniSplatting; HierarchicalCompression; RadSplat; FLoD). Scaffold-GS (ScaffoldGS) uses latent vectors anchored to reference Gaussians, with an MLP generating associated Gaussians at render time. Octree-GS (OctreeGS) extends this concept to hierarchical LoD rendering, enabling real-time control over detail levels through spatial subdivision.
3 Preliminaries
Here, we briefly review 3DGS and how Gaussian hierarchies enable adaptive level-of-detail rendering.
3.1 Hierarchical 3D Gaussian Splatting
3DGS (3DGS) models a radiance field using a set of spatially distributed Gaussians, each with mean $\bm{\mu}_{i}\in\mathbb{R}^{3}$, RGB base colors $\mathbf{b}_{i}\in\mathbb{R}^{3}$ and covariance matrices $\bm{\Sigma}_{i}=\mathbf{R}_{i}\mathbf{S}_{i}\mathbf{S}_{i}^{\top}\mathbf{R}_{i}^{\top}$, which are parameterized via a diagonal scaling matrix $\mathbf{S}_{i}=\text{diag}(s^{1}_{i},s^{2}_{i},s^{3}_{i})$ and an orthonormal rotation matrix $\mathbf{R}_{i}$. Each Gaussian also stores an opacity $\sigma_{i}$ and view-dependent color, modeled using spherical harmonics (SH) coefficients $\mathbf{f}_{i}^{d}$. The SH degree $d$ controls expressiveness, with each Gaussian requiring $\sum_{j=1}^{d}3\cdot(2j+1)$ parameters. For rendering, all $N$ Gaussians are sorted by distance to the camera and a discrete approximation of the volume rendering equation is evaluated for every pixel $\mathbf{x}$ with corresponding view direction $\mathbf{v}$:
$\mathbf{C}(\mathbf{x})=\sum_{i=1}^{N}\mathbf{c}_{i}(\mathbf{v})\alpha_{i}(\mathbf{x})\prod_{j=1}^{i-1}(1-\alpha_{j}(\mathbf{x})),$ (1)
@@ 3.1 Hierarchical 3D Gaussian Splatting
$\alpha_{j}(\mathbf{x})=\sigma_{j}e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu^{\prime}}_{j})\bm{\Sigma^{\prime}}(\mathbf{x}-\bm{\mu^{\prime}}_{j})^{T}}.$ (2)
@@
Here $\bm{\mu^{\prime}}$ and $\bm{\Sigma^{\prime}}$ denote the projected 2D mean and covariance on the image plane, obtained by applying an affine approximation of the projective transform (zwicker2001EWA).
3DGS Memory
@@
Standard 3DGS pipelines store the full set of per-Gaussian attributes, training images, and optimizer state in GPU memory (VRAM); see Figure A.1 for a detailed per-Gaussian breakdown. Additional temporary allocations occur during forward and backward passes (e.g., for sorting and gradient accumulation). This overhead varies with the scale of the Gaussians and effectiveness of culling strategies, but can be roughly upper-bounded by around 800 bytes per Gaussian in practice. This limits typical training to roughly 500 000500\,000 Gaussians per GB of GPU memory, imposing strict constraints on the detail and extent of reconstructions.
Gaussian hierarchies
@@
as introduced in H-3DGS (Hierarchical), recursively merge nearby Gaussians into a tree, where each non-leaf node approximates its children, and leaves correspond to the original Gaussians. A cut is defined by a condition $c_{\text{hier}}(i,\text{cam})$ evaluated in a breadth-first search (BFS). If a node satisfies the cut condition, it is added to the cut set and its children are skipped; otherwise, the BFS continues. A proper cut set includes no parents or children of any included node and thus provides a view-adaptive LoD representation.
The cut condition used in Hierarchical is a simple cutoff to the camera distance:
$c_{\text{hier}}(i,\text{cam})=\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i),\quad m_{d}(i)=\frac{T}{\max_{j}s_{i}^{j}},$ (3)
@@
where $\mathbf{p}_{cam}$ is the camera position, $T$ is a global LoD threshold, and $m_{d}(i)$ is the minimum acceptable distance for viewing Gaussian $i$. The BFS ensures that if $i$ is in the cut set, parent($i$) failed the condition: $m_{d}(\text{parent}(i))>\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
4 Method